# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v4

@@ II.2 Boundary zero modes and Toeplitz braiding: A concrete example
Zero modes of $K$-matrices have the following properties: for a finite size $K$-matrix, the zero modes have finite values instead of being zero. As the size of the system, i.e., the size of the $K$-matrix, increases, the eigenvalues decrease to zero exponentially. Therefore, for a finite size $K$-matrix, the inverse $K^{-1}$ is always well-defined. In the following discussion, we mainly focus on the asymptotic behavior of $K^{-1}$ as the number of layers $N\rightarrow\infty$. Thus, practically, the inverse of $K$ is calculated before taking the limit of the system size [^4] In this paper, system size refers to the dimension of a finite-size block-tridiagonal Toeplitz $K$-matrix. $2N\rightarrow\infty$.
To gain insights, we focus on a simple example, where the Lagrangian is given by Eq. (1), and the $K$-matrix (denoted as $K_{1}$-matrix hereafter) is given by Eq. (2). The $A$ and $B$ blocks read:
@@
where $m,l\in\mathbb{Z}$. Denote the $K$-matrix of size $2N$ with block $A$, $B$ given by Eq. (4) as $K_{1}^{2N}$, which is, in its mathematical form, equivalent to the Hamiltonian of Su-Schrieffer-Heeger model. The inverse $(K_{1}^{2N})^{-1}$ reads [43]
@@
It is noteworthy that in some cases, the nonzero upper-right and lower-left elements may not result in nontrivial Toeplitz braiding. For example, if $l$ is divisible by $m~{}\mathrm{rad}(m)$, the upper-right and lower-left elements in Eq. (5) are integers, implying that excitations residing at distinct $z$-boundaries cannot interact by braiding. Indeed, the presence of boundary zero modes is a necessary but not sufficient condition for the appearance of Toeplitz braiding. We have demonstrated that the appearance of boundary zero modes in $K$-matrices will result in non-zero elements in the upper-right and lower-left regions of $K^{-1}$. However, the elements in these regions are not guaranteed to be non-integers. The $K_{1}$-type iCS field theory with $m=2$, $l=4$ is a typical example. The $K$-matrix has boundary zero modes, but nontrivial braiding statistics is completely absent except for the topological excitations residing at the same layer, as shown in Fig. 4(a). The presence of non-integers in the upper-right and lower-left regions of $K^{-1}$ is necessary for the occurrence of Toeplitz braiding. Therefore, we propose the following condition for the existence of Toeplitz braiding: If boundary zero modes of $K$ along with non-integer upper-right and lower-left regions of $K^{-1}$ exist, then Toeplitz braiding occurs, which is also illustrated in Fig. 4(b).
III Sorting $K$-matrices with boundary zero modes
III.1 General approach to sorting $K$-matrices with boundary zero modes
Our investigation into a concrete example of iCS field theories with Toeplitz braiding in Sec. II raises an intriguing question: how to identify and classify iCS field theories with Toeplitz braiding? As illustrated in Fig. 4(b), the appearance of Toeplitz braiding demands both boundary zero modes of $K$ and non-integer upper-right and lower-left elements in $K^{-1}$. In this section, we present a systematic framework for determining the presence of boundary zero modes in a given $K$-matrix. As shown in Table 1, we identify two different types of $K$-matrices with boundary zero modes as illustrative examples. The Type-I $K$-matrix possesses two linearly independent boundary zero modes, whereas the Type-II $K$-matrix exhibit only one boundary zero mode. The corresponding iCS field theories, are dubbed as Type-I and Type-II iCS field theories, respectively. The aforementioned example, an iCS field theory with an Su-Schrieffer-Heeger (SSH)-type $K$-matrix, represents a typical Type-I iCS field theory.
@@ III.1 General approach to sorting $K$-matrices with boundary zero modes
Since we are discussing boundary zero modes, the $\lambda\rightarrow 0$ limit is taken at first. In terms of the Toeplitz $K$-matrices in the form of Eq. (2), the entries of a trial solution $\mathbf{u}$ to boundary zero mode are required to satisfy the recurrence relation
@@
$\beta_{j}$ together with the normalization condition $(\mathbf{w}^{j})^{\dagger}\mathbf{w}^{j}=1$ settles down the entries of $\mathbf{w}^{j}$. Furthermore, we require that $\mathbf{u}$ localizes at one $z$-boundary, i.e., all the $\{\beta_{j}\}$ in one trial solution $\mathbf{u}$ satisfies either $|\beta_{j}|>1$ or $|\beta_{j}|<1$. If a trial solution $\mathbf{u}$ is a linear combination of boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|<1$, the boundary condition Eq. (16b) is automatically satisfied in the thermodynamical limit $2N\rightarrow\infty$. Hence our task is to verify whether a linear combination of boundary bases (given by Eq. (17)) exists, matching the lower boundary condition Eq. (16a). Likewise, for boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|>1$, our task is to verify whether a linear combination of boundary bases matching the upper boundary condition Eq. (16b) exists. If $\mathbf{u}$ satisfies the lower boundary condition Eq. (16a), it is referred to as lower boundary zero mode. If $\mathbf{u}$ satisfies the upper boundary condition Eq. (16b), it is referred to as upper boundary zero mode. Following this framework, we can systematically determine whether a given $K$-matrix possesses boundary zero modes.
In the preceding discussion, we made the assumption that the $K$-matrix is infinite-dimensional. However, in the study of Toeplitz braiding, we calculate the inverse of a $K$-matrix before taking the thermodynamic limit $2N\rightarrow\infty$. Therefore, it becomes essential to incorporate the discussion on the boundary zero modes and the corresponding exponentially-suppressed eigenvalues of a finite-size $K$-matrix. If an infinite-dimensional $K$-matrix possesses two boundary zero modes, $\mathbf{u}^{1}$ and $\mathbf{u}^{2}$, satisfying distinct boundary conditions Eq. (16a) and Eq. (16b), respectively, the actual boundary zero modes for a finite-size $K$-matrix are linear combinations of $\mathbf{u}^{1}$ and $\mathbf{u}^{2}$, which can be obtained by diagonalizing the matrix
@@
If an infinite-dimensional $K$-matrix possesses only one boundary zero mode $\mathbf{u}$ meeting either Eq. (16a) or Eq. (16b), it is in fact not an actual boundary zero mode for a finite-size $K$-matrix due to the violation to the other boundary condition. The violation will provide a minor correction to the entries of $\mathbf{u}$, contributing a small but nonzero eigenvalue $\lambda$, which enables the calculation of the inverse of $K$. In Sec. IV, we will demonstrate that $M=\frac{1}{\lambda}\mathbf{u}\mathbf{u}^{\dagger}$ also captures the braiding statistics of topological excitations residing at distinct $z$-boundaries, in analogy to the cases of two boundary zero modes.
In the subsequent subsections, we provide a detailed discussion on sorting the Type-I and Type-II Toeplitz $K$-matrices. For a $K$-matrix as shown in Eq. (2), we follow a two-step procedure. First, we solve for all the boundary bases, then we verify whether a linear superposition of these boundary bases exists, as expressed in Eq. (17), satisfying either the upper boundary condition Eq. (16a) or lower boundary condition Eq. (16b). Therefore, for a given (twisted) topological order described by $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$, we can identify the specific Chern-Simons coupling associated with the block $B$ in the $K$-matrix that enables Toeplitz braiding. We organize our analysis based on whether the determinant of matrix $B$ is equal to zero or not. $\det B$ plays a crucial role in determining the order of Eq. (21), which affects the # of boundary bases.
@@ III.2 Discussion on block $B$
from which we know that $\beta_{1}\beta_{2}=1$. For the sake of convenience, we set $|\beta_{1}|<1$, $|\beta_{2}|>1$. The bulk recurrence relation Eq. (19) together with normalization condition gives the solution to $\begin{pmatrix}w^{j,a}&w^{j,b}\end{pmatrix}^{T}$, $j=1,2$. If $\mathbf{u}^{1}=\mathbf{w}^{1}$ is a legitimate boundary zero mode, it satisfies the boundary condition Eq. (16a). Detailed calculation in Appendix A.1 shows that there exists one boundary zero mode satisfying the lower boundary condition Eq. (16a) if
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\\ 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{2}=l_{4}=0,\\ |m|<|l_{3}|.\end{array}\right.$ (36)
Likewise, there exists boundary zero mode $\mathbf{u}^{2}=\mathbf{w}^{2}$ satisfying upper boundary condition Eq. (16b) if
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\\ 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|.\\ \end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{3}=l_{4}=0,\\ |m|<|l_{2}|.\end{array}\right.$ (44)
Combining the condition (36) and (44), we conclude that if
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{1}=l_{3}=l_{4}=0,\\ |m|<|l_{2}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{1}=l_{2}=l_{4}=0,\\ |m|<|l_{3}|,\end{array}\right.$ (51)
the $K$-matrix exhibits two boundary zero modes. Otherwise, if $\det B=0$, and the elements of Eq. (2) satisfies only one of the two conditions (36) and (44), the $K$-matrix possesses only 1 boundary zero mode.
@@
which enables a thorough discussion. Detailed analysis are shown in Appendix A.2. We discover that if
$\left\{\begin{array}[]{l}l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\\ l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0,\end{array}\right.$ (57)
the $K$-matrix possesses one boundary zero mode. More specifically, for the case $|l_{2}|<|l_{3}|$, the boundary zero mode satisfies the lower boundary condition Eq. (94); for the case $|l_{2}|>|l_{3}|$, the boundary zero mode satisfies the lower boundary condition Eq. (95). If
$\left\{\begin{array}[]{l}l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\\ n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0,\end{array}\right.$ (62)
the $K$-matrix possesses two boundary zero modes.

### v5

@@ II.2 Boundary zero modes and Toeplitz braiding: A concrete example
Zero modes of $K$ matrices have the following properties: for a finite size $K$ matrix, the zero modes have finite values instead of being zero. As the size of the system, i.e., the size of the $K$ matrix, increases, the eigenvalues decrease to zero exponentially. Therefore, for a finite size $K$ matrix, the inverse $K^{-1}$ is always well-defined. In the following discussion, we mainly focus on the asymptotic behavior of $K^{-1}$ as the number of layers $N\rightarrow\infty$. Thus, practically, the inverse of $K$ is calculated before taking the limit of the system size [^4] In this paper, system size refers to the dimension of a finite-size block-tridiagonal Toeplitz $K$ matrix. $2N\rightarrow\infty$.
To gain insights, we focus on a simple example, where the Lagrangian is given by Eq. (1), and the $K$ matrix (denoted as $K_{1}$-matrix hereafter) is given by Eq. (2). The $A$ and $B$ blocks read:
@@
where $m,l\in\mathbb{Z}$. Denote the $K$ matrix of size $2N$ with block $A$, $B$ given by Eq. (4) as $K_{1}^{2N}$, which is, in its mathematical form, equivalent to the Hamiltonian of Su-Schrieffer-Heeger model. The inverse $(K_{1}^{2N})^{-1}$ reads [44]
@@
It is noteworthy that in some cases, the nonzero upper-right and lower-left elements may not result in nontrivial Toeplitz braiding. For example, if $l$ is divisible by $m~{}\mathrm{rad}(m)$, the upper-right and lower-left elements in Eq. (5) are integers, implying that excitations residing at distinct $z$-boundaries cannot interact by braiding. Indeed, the presence of boundary zero modes is a necessary but not sufficient condition for the appearance of Toeplitz braiding. We have demonstrated that the appearance of boundary zero modes in $K$ matrices will result in non-zero elements in the upper-right and lower-left regions of $K^{-1}$. However, the elements in these regions are not guaranteed to be non-integers. The $K_{1}$-type iCS field theory with $m=2$, $l=4$ is a typical example. The $K$ matrix has boundary zero modes, but nontrivial braiding statistics is completely absent except for the topological excitations residing at the same layer, as shown in Fig. 4(a). The presence of non-integers in the upper-right and lower-left regions of $K^{-1}$ is necessary for the occurrence of Toeplitz braiding. Therefore, we propose the following condition for the existence of Toeplitz braiding: If boundary zero modes of $K$ along with non-integer upper-right and lower-left regions of $K^{-1}$ exist, then Toeplitz braiding occurs, which is also illustrated in Fig. 4(b).
III Sorting $K$ matrices with boundary zero modes
III.1 General approach to sorting $K$ matrices with boundary zero modes
Our investigation into a concrete example of iCS field theories with Toeplitz braiding in Sec. II raises an intriguing question: how to identify and classify iCS field theories with Toeplitz braiding? As illustrated in Fig. 4(b), the appearance of Toeplitz braiding demands both boundary zero modes of $K$ and non-integer upper-right and lower-left elements in $K^{-1}$. In this section, we present a systematic framework for determining the presence of boundary zero modes in a given $K$ matrix. As shown in Table 1, we identify two different types of $K$ matrices with boundary zero modes as illustrative examples. The Type-I $K$ matrix possesses two linearly independent boundary zero modes, whereas the Type-II $K$ matrix exhibit only one boundary zero mode. The corresponding iCS field theories, are dubbed as Type-I and Type-II iCS field theories, respectively. The aforementioned example, an iCS field theory with an Su-Schrieffer-Heeger (SSH)-type $K$-matrix, represents a typical Type-I iCS field theory.
@@ III.1 General approach to sorting $K$ matrices with boundary zero modes
Since we are discussing boundary zero modes, the $\lambda\rightarrow 0$ limit is taken at first. In terms of the Toeplitz $K$ matrices in the form of Eq. (2), the entries of a trial solution $\mathbf{u}$ to boundary zero mode are required to satisfy the recurrence relation
@@
$\beta_{j}$ together with the normalization condition $(\mathbf{w}^{j})^{\dagger}\mathbf{w}^{j}=1$ settles down the entries of $\mathbf{w}^{j}$. Furthermore, we require that $\mathbf{u}$ localizes at one $z$-boundary, i.e., all the $\{\beta_{j}\}$ in one trial solution $\mathbf{u}$ satisfies either $|\beta_{j}|>1$ or $|\beta_{j}|<1$. If a trial solution $\mathbf{u}$ is a linear combination of boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|<1$, the boundary condition Eq. (16b) is automatically satisfied in the thermodynamical limit $2N\rightarrow\infty$. Hence our task is to verify whether a linear combination of boundary bases (given by Eq. (17)) exists, matching the lower boundary condition Eq. (16a). Likewise, for boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|>1$, our task is to verify whether a linear combination of boundary bases matching the upper boundary condition Eq. (16b) exists. If $\mathbf{u}$ satisfies the lower boundary condition Eq. (16a), it is referred to as lower boundary zero mode. If $\mathbf{u}$ satisfies the upper boundary condition Eq. (16b), it is referred to as upper boundary zero mode. Following this framework, we can systematically determine whether a given $K$ matrix possesses boundary zero modes.
In the preceding discussion, we made the assumption that the $K$ matrix is infinite-dimensional. However, in the study of Toeplitz braiding, we calculate the inverse of a $K$ matrix before taking the thermodynamic limit $2N\rightarrow\infty$. Therefore, it becomes essential to incorporate the discussion on the boundary zero modes and the corresponding exponentially-suppressed eigenvalues of a finite-size $K$ matrix. If an infinite-dimensional $K$ matrix possesses two boundary zero modes, $\mathbf{u}^{1}$ and $\mathbf{u}^{2}$, satisfying distinct boundary conditions Eq. (16a) and Eq. (16b), respectively, the actual boundary zero modes for a finite-size $K$ matrix are linear combinations of $\mathbf{u}^{1}$ and $\mathbf{u}^{2}$, which can be obtained by diagonalizing the matrix
@@
If an infinite-dimensional $K$ matrix possesses only one boundary zero mode $\mathbf{u}$ meeting either Eq. (16a) or Eq. (16b), it is in fact not an actual boundary zero mode for a finite-size $K$ matrix due to the violation to the other boundary condition. The violation will provide a minor correction to the entries of $\mathbf{u}$, contributing a small but nonzero eigenvalue $\lambda$, which enables the calculation of the inverse of $K$. In Sec. IV, we will demonstrate that $M=\frac{1}{\lambda}\mathbf{u}\mathbf{u}^{\dagger}$ also captures the braiding statistics of topological excitations residing at distinct $z$-boundaries, in analogy to the cases of two boundary zero modes.
In the subsequent subsections, we provide a detailed discussion on sorting the Type-I and Type-II Toeplitz $K$ matrices. For a $K$ matrix as shown in Eq. (2), we follow a two-step procedure. First, we solve for all the boundary bases, then we verify whether a linear superposition of these boundary bases exists, as expressed in Eq. (17), satisfying either the upper boundary condition Eq. (16a) or lower boundary condition Eq. (16b). Therefore, for a given (twisted) topological order described by $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$, we can identify the specific Chern-Simons coupling associated with the block $B$ in the $K$ matrix that enables Toeplitz braiding. We organize our analysis based on whether the determinant of matrix $B$ is equal to zero or not. $\det B$ plays a crucial role in determining the order of Eq. (21), which affects the # of boundary bases.
@@ III.2 Discussion on block $B$
from which we know that $\beta_{1}\beta_{2}=1$. For convenience, we assume $|\beta_{1}|<1$, $|\beta_{2}|>1$. The bulk recurrence relation Eq. (19) together with normalization condition gives the solution to $\begin{pmatrix}w^{j,a}&w^{j,b}\end{pmatrix}^{T}$, $j=1,2$. If $\mathbf{u}^{1}=\mathbf{w}^{1}$ is a legitimate boundary zero mode, it satisfies the boundary condition Eq. (16a). Detailed calculation in Appendix B shows that there exists a boundary zero mode satisfying the lower boundary condition Eq. (16a) if the condition (39) is satisfied.
$\mathbb{A}_{1}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\text{ or }\mathbb{A}_{2}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{2}=l_{4}=0,\\ |m|<|l_{3}|,\end{array}\right.\text{ or }\mathbb{A}_{3}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ nl_{2}=2l_{1}m,\\ l_{3}=l_{4}=0,\\ |m|<|l_{2}|.\end{array}\right.$ (39)
Likewise, there exists a boundary zero mode $\mathbf{u}^{2}=\mathbf{w}^{2}$ satisfying the upper boundary condition Eq. (16b) if the boundary condition (50) is satisfied.
$\mathbb{B}_{1}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\quad l_{2}\neq l_{3},\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\\ \end{array}\right.\text{ or }\mathbb{B}_{2}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{3}=l_{4}=0,\\ |m|<|l_{2}|,\end{array}\right.\text{ or }\mathbb{B}_{3}:\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ nl_{3}=2l_{1}m,\\ l_{2}=l_{4}=0,\\ |m|<|l_{3}|.\end{array}\right.$ (50)
Here Eq. (39) represents the union $\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3}$ of the sets $\mathbb{A}_{1,2,3}$. Likewise, Eq. (50) represents the union $\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3}$. Intersecting the above conditions Eq. (39) and Eq. (50), i.e. the intersection $(\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3})\cap(\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3})$, we should have 9 sets of conditions $\mathbb{A}_{1,2,3}\cap\mathbb{B}_{1,2,3}$ to ensure that the $K$ matrix has two boundary zero modes. With further analysis, only $\mathbb{A}_{3}\cap\mathbb{B}_{2}$ and $\mathbb{A}_{2}\cap\mathbb{B}_{3}$ are not empty set, where the remaining sets are empty set. Therefore, we conclude that if $\det B=0$, and the elements of the $K$ matrix satisfies
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{2}\neq 0,\\ l_{3}=l_{4}=0,\\ nl_{2}=2l_{1}m,\\ |m|<|l_{2}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{2}=l_{4}=0,\\ nl_{3}=2l_{1}m,\\ |m|<|l_{3}|,\end{array}\right.$ (59)
the $K$ matrix exhibits two boundary zero modes. Otherwise, if $\det B=0$, and the elements of Eq. (2) satisfies only one of the two conditions Eq. (39) and Eq. (50), the $K$ matrix possesses only one boundary zero mode. Furthermore, the conditions for only one boundary zero mode under $\det B=0$ are summarized as follows:
$\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\\ 2l_{3}m=l_{4}n\ \text{ or }\ 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\text{ or }\left\{\begin{array}[]{l}m\neq 0,\quad l_{3}\neq 0,\\ l_{4}=l_{2}=0,\\ nl_{3}\neq 2l_{1}m\\ |m|<|l_{3}|,\end{array}\right.$ (68)
[…]
Eq. (73) is obtained by the set operation $(\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3}\cup\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3})\backslash[(\mathbb{A}_{1}\cup\mathbb{A}_{2}\cup\mathbb{A}_{3})\cap(\mathbb{B}_{1}\cup\mathbb{B}_{2}\cup\mathbb{B}_{3})]$. More specifically, if $\left\{\begin{array}[]{l}l_{4}=0\\ |l_{2}|<|l_{3}|\end{array}\right.\text{ or }\left\{\begin{array}[]{l}l_{4}\neq 0\\ 2l_{3}m=l_{4}n\end{array}\right.$, the boundary zero mode is a lower boundary zero mode, satisfying the boundary condition Eq. (16a). If $\left\{\begin{array}[]{l}l_{4}=0\\ |l_{3}|<|l_{2}|\end{array}\right.\text{ or }\left\{\begin{array}[]{l}l_{4}\neq 0\\ 2l_{2}m=l_{4}n\end{array}\right.$, the boundary zero mode is an upper boundary zero mode, satisfying the boundary condition Eq. (16b).
@@
which enables a thorough discussion. Detailed calculation in Appendix C shows that there exists a boundary zero mode satisfying the lower boundary condition Eq. (16a) if the following conditions (82) are satisfied.
$\mathbb{C}_{1}:\ \left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\\ \end{array}\right.\quad\text{or}$ (78)
$\mathbb{C}_{2}:\ \left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\end{array}\right.$ (82)
Likewise, there exists one boundary zero mode satisfying the upper boundary condition Eq. (16b) if the following conditions (90) are satisfied.
$\mathbb{D}_{1}:\ \left\{\begin{array}[]{l}l_{4}=0,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\end{array}\right.\quad\text{or}$ (86)
$\mathbb{D}_{2}:\ \left\{\begin{array}[]{l}l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ m\neq 0,\quad l_{2}l_{3}\neq 0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\\ \end{array}\right.$ (90)
Here Eq. (82) and Eq. (90) represents the union $\mathbb{C}_{1}\cup\mathbb{C}_{2}$ and $\mathbb{D}_{1}\cup\mathbb{D}_{2}$, respectively. Set operation $(\mathbb{C}_{1}\cup\mathbb{C}_{2})\cap(\mathbb{D}_{1}\cup\mathbb{D}_{2})$ yields the following condition that ensures the $K$ matrix (Eq. (2)) possessing two boundary zero modes:
$\left\{\begin{array}[]{l}l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\\ n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0.\end{array}\right.$ (95)
Furthermore, set operation $(\mathbb{C}_{1}\cup\mathbb{C}_{2}\cup\mathbb{D}_{1}\cup\mathbb{D}_{2})\backslash[(\mathbb{C}_{1}\cup\mathbb{C}_{2})\cap(\mathbb{D}_{1}\cup\mathbb{D}_{2})]$ yields the following condition for the appearance of exactly one boundary zero mode:
$\left\{\begin{array}[]{l}l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\\ l_{2}\neq l_{3},\quad l_{2}l_{3}\neq 0.\end{array}\right.$ (100)
More specifically, for the case $|l_{2}|<|l_{3}|$, the boundary zero mode satisfies the lower boundary condition Eq. (16a); for the case $|l_{2}|>|l_{3}|$, the boundary zero mode satisfies the upper boundary condition Eq. (16b).